OpenStudy (anonymous):

3x-2=10 have to identify as horizontal or vertical n' graph.

OpenStudy (mertsj):

Solve it for x.

OpenStudy (mertsj):

@Jlynn388 What did you get?

OpenStudy (anonymous):

don't you get all the numbers on 1 side so you get y=mx+b ?

OpenStudy (anonymous):

I did 3x-2=10 added 2 to both sides so 3x=12 then divided both by 3 so 12/3 x= 4

OpenStudy (mertsj):

Yes. Now, get some ordered pairs on that line. Notice that x has to be 4 according to the equation and the equation says nothing about y so y can be anything you choose.

OpenStudy (mertsj):

What ordered pairs did you write?

OpenStudy (anonymous):

4,3 4,-1

OpenStudy (mertsj):

Graph them. And draw the line.

OpenStudy (mertsj):

Is the line horizontal or vertical?

OpenStudy (mertsj):